ganeshie8 (ganeshie8):

show that every odd prime divisor of \(3^{(2^n)}-1\) is of form \(4k+1\)

ganeshie8 (ganeshie8):

I have a really nice proof, but it uses many luxury theorems. Looking for an elementary proof if it is available... :)

ganeshie8 (ganeshie8):

Haha! I see what you have discovered... It seems the sequence is spitting out a "new" prime in the prime factorization of each subsequent term...

ganeshie8 (ganeshie8):

It doesn't break at 7 : 2^3, 2^4×5, 2^5×5×41, 2^6×5×17×41×193, 2^7×5×17×41×193×21523361, 2^8×5×17×41×193×21523361×926510094425921 2^9×5×17×41×193×21523361×926510094425921×1716841910146256242328924544641

Parth (parthkohli):

Oh, never mind.

ganeshie8 (ganeshie8):

I have checked first 11 terms, each has a new prime

Parth (parthkohli):

Yeah, makes sense. And that prime is of course, of the form, \(4k+1\). I was actually thinking of proving \(3^{2^n}+1\) has the same property as mentioned in the question, so that if we multiply this and \(3^{2^n}-1\), we inductively have the next term.

ganeshie8 (ganeshie8):

Looks you're onto something !

ganeshie8 (ganeshie8):

Let \(F_n = 3^{2^n}-1\) basically you're saying \(F_{n+1} =(F_n+2)F_n \)

OpenStudy (kainui):

I found this quite strange too, that it seems like we're proving that \(3^{2^n} \pm 1\) all have odd prime divisors of the form 4k+1 which seems like a kind of big deal.

Parth (parthkohli):

Exactly. I brought it into that form and all, but couldn't follow through.

ganeshie8 (ganeshie8):

Yeah it is amazing there are no online references about this I'm not sure where @imqwerty found this problem...

imqwerty (imqwerty):

(: try this i think u guys can do it i'll post the solution soon

ganeshie8 (ganeshie8):

@ParthKohli that factorization also gives a trivial result : \(p\mid F_n \implies p\mid F_{n+1}\) that means \(F_k\mid F_{n}\) for all \(k\le n\)

Parth (parthkohli):

yeah but it's the \(F_n + 2\) term that's troubling

Parth (parthkohli):

@imqwerty were you able to solve this

OpenStudy (kainui):

I don't think what I'm about to say will be particularly important to this problem, however it seems interesting. I noticed a peculiar thing with exponents m > 2 (where have I heard this kind of thing before...?) that took me off track, this inequality here: \[\sum_{k=0}^{m^n-1}a^k > \prod_{k=0}^{n-1} (a^{m^k}+1)\] However it is an equality when m = 2 because since the right side will then have every permutation of sums of powers of 2 (whew) which in binary works out to be all the numbers less than 1's for n-1 times... You can't do this in base 3 or higher, since we need to multiply a power of 3 by 2 to get more than just the 0 and 1 digits if this makes sense.... lol.

imqwerty (imqwerty):

yes pk :) my method is lil long m trying to shorten it i think u guys can get it too

Parth (parthkohli):

ok i'm going to go to a corner and cry myself to sleep